OpenStudy (anonymous):

Differentiate (x+3√x)' show work using first' last first last' method

OpenStudy (anonymous):

the what?

Parth (parthkohli):

I'd use the power rule instead.

OpenStudy (cwrw238):

i've never heard of that!!

OpenStudy (anonymous):

its ok how would you do it

OpenStudy (anonymous):

derivative of \(x\) is 1 derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\)

Parth (parthkohli):

If it's a multiple choice, then you may screw the method given in your question. Embrace the power rule.

OpenStudy (anonymous):

someones math teacher must have made this up. it is not standard. probably thought they were being clever

OpenStudy (anonymous):

its not multiple choice

OpenStudy (cwrw238):

first last first last - it sounds biblical!!

OpenStudy (anonymous):

the last shall be first the greek shall inherit the earth the answer is \(1+\frac{3}{2\sqrt{x}}\)

Parth (parthkohli):

The Dark Knight shall rise in fiction.

OpenStudy (cwrw238):

lol

Parth (parthkohli):

Cats shall be posted as profile pics Bikes shall be green